# RIS-aided Wireless Communication with Movable Elements: Geometry Impact on Performance

### v1

@@ II-B Topology of MA-RIS illumination area
Let us consider the system model shown in Fig. 1. All the elements are moving horizontally, with the same constant speed $v$ from the RIS’s left panel edge toward the right edge. By fitting the illumination area into an elliptic shape, we can derive the illumination area that affects the UE’s SNR. In order to capture the most important feature in the context of MA-RIS – the AoA and AoD information related with the movement of MA – we followed the cosine model in [14, 5], and for the $n^{th}$ element of MA-RIS, its radiation pattern for the incident beam from BS is described using cosine model which is given as $U_{t}(\varphi_{n}^{r})=\mathrm{cos}^{2}\left(\frac{\pi(\varphi_{n}^{r}-\varphi_{r})}{2\phi_{0}}\right)$, where the HPBW of the transmitted beam is denoted as $\phi_{0}$, with the angular information $\varphi_{n}^{r}=\mathrm{arctan}\frac{x_{n}^{r}}{y_{r}}$, and with $\varphi_{r}=\mathrm{arctan}\frac{x_{r}}{y_{r}}$ representing the azimuth AoA for MA-RIS center. The radiation pattern of the $n^{th}$ element of MA-RIS corresponding to the incident and reflect beam is separately defined in a similar way as in [14], as $W_{r}(\varphi_{n}^{r})=\mathrm{cos}^{3}(\varphi_{n}^{r})$ and $W_{t}(\varphi_{n}^{t})=\mathrm{cos}^{3}(\varphi_{n}^{t})$ with $\varphi_{n}^{t}=\mathrm{arctan}\frac{x^{u}_{k}-x^{r}_{n}}{y^{u}_{k}-y^{r}_{n}}$ representing the azimuth AoD from $n^{th}$ MA-RIS element to $k^{th}$ UE. Therefore, we get the received power at the UE as
$P_{r}=P_{t}\frac{\lambda^{2}d_{x}d_{y}}{64\pi^{3}}\left|\sum_{i=1}^{N_{eff}}\frac{\sqrt{U_{r}(\varphi_{n}^{r})W_{r}(\varphi_{n}^{r})W_{t}(\varphi_{n}^{t})}}{d_{n,1}d_{n,2}}\right|^{2}$ (14)
where $P_{t}$ is the transmit power, $d_{x}\times d_{y}$ represents the size of a flat MA-RIS element, $d_{n,1}=\left\|\textbf{b}_{m}-\textbf{r}_{n}\right\|$ and $d_{n,2}=\left\|\textbf{r}_{n}-\textbf{u}_{k}\right\|$ denote the 3D distance between BS to $n^{th}$ MA-RIS element and the 3D distance between the $n^{th}$ MA-RIS element to $k^{th}$ UE, respectively.
[…]
@@ II-C Outage Probability
We use the same method to evaluate the performance of outage probability as [5], defined as the probability that the SNR is smaller than a given threshold $\gamma_{th}$,
$P_{out}(\gamma_{th})=\mathrm{Prob}(\gamma_{max}\leq\gamma_{th})=F_{\gamma_{max}}(\gamma_{th})$ (15)
where $F_{\gamma_{max}}(\cdot)$ denotes the Cumulative Distribution Function (CDF) of $\gamma_{max}$, where $\gamma_{max}=\bar{\gamma}\left({\textstyle\sum_{i=1}^{N_{eff}}\eta_{i}}\right)^{2}$, where we denote $\eta_{i}=\alpha_{i}\beta_{i}$. With the method of moment matching [15], we can obtain $F_{\gamma_{max}}=\frac{\Gamma\left(\Lambda N_{eff},\sqrt{\frac{\gamma_{th}}{\bigtriangleup^{2}\bar{\gamma}}}\right)}{\Gamma\left(\Lambda N_{eff}\right)}$, where $\Gamma\left(\cdot,\cdot\right)$ is the incomplete Gamma function, while $\Gamma\left(\cdot\right)$ is the Gamma function.
@@ III Performance Analysis
We analyse the performance in near field and assume that the elements within the MA-RIS configuration shift horizontally, transitioning from the panel’s left edge to its right edge. The positions of BS and UE are set as $(x_{b},y_{b},z_{b})=(0,0,3)$m and $(x^{u}_{k},y^{u}_{k},z^{u}_{k})=(10,0,0)$m, the carrier frequency is $f=5.25$ GHz, the noise variance is $\sigma^{2}=-45$ dBm. The RIS panel length is 1m, and the number of elements is $N=21$. The antenna movement velocity is $v=0.1$ m/s, with the traversal time thus amounting to $t=10$s .
Figure 2: Outage probability performance comparison. (a) $P_{out}$ versus $P_{t}$. (b) $P_{out}$ versus elements number.
@@ A Transmit Power and Number of Elements : Effect on Outage probability
As shown in Fig.2LABEL:sub@OutvsPt, the outage probability decreases with an increase in transmit power for all configurations. In 2D scenario with 21 elements and -20 dBm transmit power, the MA-RIS achieves around 24% better performance than FPA-RIS in terms of outage probability. In a 2D scenario, the outage probability sharply decreases with increasing transmit power due to better control over signal propagation and interference mitigation. This is because the additional dimension allows for more precise optimization of signal paths, resulting in improved channel conditions. The impact of number of elements on outage probability $P_{out}$ is shown in Fig.2LABEL:sub@OutvsN, with the transmit power $P_{t}$ fixed at -20 dBm. Despite being limited by the single-path condition, there is no significant difference between the 1D and 2D configurations for both the MA-RIS and FPA-RIS settings. In particular, the results show that MA-RIS uses fewer elements than FPA-RIS to achieve the same level of performance, for example for an outage probability of $10^{-70}$, MA-RIS uses 25% fewer elements than FPA-RIS.
These promising results are the consequence of MA-enabled dynamic reconfiguration of antenna elements, allowing for adaptive adjustment of the antenna pattern and phase distribution. This flexibility can be leveraged to optimize the signal path and mitigate multipath fading, leading to enhanced communication reliability and lower outage probabilities compared to FPA-RIS, which have a limited capability to adapt to changing channel conditions.
@@ B Effective Number of MA-RIS
We set both the MA-RIS and FPA-RIS panels are depicted moving along the y-axis, denoted as $y_{s}$ in Fig.3. They traverse from an initial position of 1m to the maximum height of 15m in increments of 2 meters. In order to compare the performance of FPA-RIS and MA-RIS over the same overall panel length, the antenna spacing for FPA-RIS is $d_{sf}=\lambda/2$, while the antenna spacing for MA-RIS is $(vt-(N-1)d_{sf})/(N-1)$. We can observe that for both 1D and 2D configurations for the MA-RIS, the effective antenna number $N_{eff}$ reaches its maximum value for lower values of $y_{s}$ as compared to FPA-RIS. This indicates that the ability to adjust the position of the RIS panel allows for optimizing the reflective path and thus the effective utilization of more antenna elements. We also notice that in 2D configurations, the RIS can quickly reach its maximum $\textit{N}_{eff}$ as it begins to move from the initial position along the y-axis. This implies that, the 2D RIS has a higher initial capacity for effective signal reflection due to its expanded interaction field, reaching peak performance with minimal movement. Instead, for 1D configurations, the effectiveness of the RIS in terms of $\textit{N}_{eff}$ increases more significantly as one moves to higher positions along the y-axis. In fact, in 1D setups the initial positions may not be optimal for capturing and redirecting signals effectively, while as the RIS moves to higher y-axis positions, it can encounter angles and positions where the linear array of antennas becomes more aligned with the direction of the incident signals, thereby increasing the effective number of antennas that can contribute to signal reflection.
Figure 3: Comparisons of effective numbers in MA-RIS and FPA-RIS versus $y_{s}$
@@ C Instantaneous SNR
In Fig.4, we compared instantaneous SNR versus the transmit power $P_{t}$. We can see for example that when the transmit power $P_{t}$ is 0 dBm, then the MA-RIS’s instantaneous SNR is about 3 dBm and 2.5 dBm higher than FPA-RIS in 1D and 2D setups, respectively. This is due to the fact that MA-RIS systems have the inherent advantage of being able to adapt their position and orientation, in response to the environment and the direction of the incoming signal. This adaptability allows them to optimize the reflection path, enhancing the signal’s strength at the receiver and thereby improving the SNR. The performance of MA-RIS and FPA-RIS in both 1D and 2D configurations are nearly the same.
We can also observe that the 2D configurations offer superior performance compared to 1D configurations, this is due to the fact that the instantaneous SNR $\gamma$ is determined by $P_{r}$, 2D has higher geometric compactness than 1D which leads to the difference of SNR, with slight difference regardless with MA-RIS and FPA-RIS settings. 2D configurations allow for more sophisticated beamforming strategies, leveraging the additional dimension to focus the energy more precisely towards the receiver. Moreover, 2D RIS configurations can accommodate a wider range of AoA and AoD values for the signal. This capability enables them to maintain high SNR levels over a variety of transmission and reception scenarios, unlike 1D RIS, which is more limited in this respect.
In 1D configurations, the FPA-RIS and MA-RIS achieves nearly the same performance. This is due to the fact that the geometric compactness leads to the difference of $y_{s}$ and the 2D RIS can achieve the highest average SNR.
Figure 4: Instantaneous SNR $\gamma$ versus transmit power $P_{t}$
@@ IV Conclusion
Our work highlights the importance of considering the following factors with regard to antenna elements: the element geometry distribution, i.e., 1D or 2D, and each element’s position in RIS panel, as well as the position of the RIS panel. Those factors are deeply connected with the angles of incidence and interaction with reflection signals, which contribute the signal strength and coverage. This could prove particularly beneficial in dynamic or large-scale environments, e.g., V2X and Massive MIMO communication systems, since the flexible adjustment of the antenna positions could provide an important additional degree of freedom to capture the geometry change, resulting in enhanced coverage, and improved signal quality. Despite our work highlight advantages of MA-RIS, the limitations inherent in our current model set the stage for our future efforts to refine and expand its capabilities. The model’s idealized assumptions, such as a fixed single-user position and a single-path condition, have resulted in a notably low outage probability. Our future work will aim to address these limitations by incorporating more realistic user scenarios and multiple path conditions, enhancing both the model’s applicability and accuracy.

### v2

@@ II-B Topology of MA-RIS illumination area
Let us consider the system model shown in Fig. 1. All the elements are moving horizontally, with the same constant speed $v$ from the RIS’s left panel edge toward the right edge. By fitting the illumination area into an elliptic shape, we can derive the illumination area that affects the Rx’s SNR. In order to capture the most important feature in the context of MA-RIS – the angle of arrival and departure information related to the movement of MA – we followed the cosine model in [15, 4], and for the $n$-th element of MA-RIS, its radiation pattern for the incident beam from Tx is then given by $U_{t}(\varphi_{n}^{r})=\mathrm{cos}^{2}\left(\frac{\pi(\varphi_{n}^{r}-\varphi_{r})}{2\phi_{0}}\right)$, where the HPBW of the transmitted beam is denoted as $\phi_{0}$, with the angular information $\varphi_{n}^{r}=\mathrm{arctan}\left(x_{n}^{r}/y_{n}^{r}\right)$, and with $\varphi_{r}=\mathrm{arctan}\frac{x_{r}}{y_{r}}$ representing the AoA for MA-RIS center. The radiation pattern of the $n$-th element of MA-RIS corresponding to the incident and reflected beam, is separately defined in a similar way as in [15], as $W_{r}(\varphi_{n}^{r})=\mathrm{cos}^{3}(\varphi_{n}^{r})$ and $W_{t}(\varphi_{n}^{t})=\mathrm{cos}^{3}(\varphi_{n}^{t})$ with $\varphi_{n}^{t}=\mathrm{arctan}\frac{x_{u}-x^{r}_{n}}{y_{u}-y^{r}_{n}}$ representing the AoD from $n$-th MA-RIS element to Rx. We deploy the near field path loss model in [11] and get the received power at the Rx as
$P_{r}=P_{t}\frac{\lambda^{2}d_{x}d_{y}}{64\pi^{3}}\left|\sum_{n=1}^{N_{eff}}\frac{\sqrt{U_{r}(\varphi_{n}^{r})W_{r}(\varphi_{n}^{r})W_{t}(\varphi_{n}^{t})}}{d_{n,1}d_{n,2}}\right|^{2}\,$, (11)
where $d_{x}\times d_{y}$ represents the size of a flat MA-RIS element, $d_{n,1}=\left\|\textbf{b}-\textbf{r}_{n}\right\|$ and $d_{n,2}=\left\|\textbf{r}_{n}-\textbf{u}\right\|$ denote the 3D distance between Tx to $n$-th MA-RIS element and the 3D distance between the $n$-th MA-RIS element to Rx, respectively.
We model the illumination area of the MA-RIS as an elliptical shape with area size as $S_{m}=\pi ab$. The semi-major and semi-minor axes of the ellipse area are calculated as $a=\frac{d_{c}\mathrm{sin}\phi_{0}}{\mathrm{sin}(\varphi_{t}+\phi_{0})}$ and $b=\frac{d_{c}\mathrm{sin}\phi_{0}}{\mathrm{cos}(\theta_{r}+\phi_{0})}$ where $d_{c}$ is the 3D distance between Tx and the center of the MA-RIS panel, and $\theta_{r}$ is the EoA from Tx to MA-RIS’s center. Therefore, we can get the effective illuminated number of the 1D and 2D MA-RIS as $N^{1D}_{eff}=\left\lceil\frac{a}{d_{s}}\right\rceil$, $N^{2D}_{eff}=\left\lceil\frac{S_{m}}{2d_{s}^{2}}\right\rceil$, where $d_{s}$ is the inter-element spacing, and $\left\lceil\cdot\right\rceil$ represents the ceiling function.
@@ II-C Outage Probability
Outage probability is a widely used metric for assessing performance in communication systems [4]. It is defined as the probability that the SNR is smaller than a given threshold $\gamma_{th}$,
$P_{out}(\gamma_{th})=\mathrm{Prob}(\gamma_{max}\leq\gamma_{th})=F_{\gamma_{max}}(\gamma_{th})\,$, (12)
where $F_{\gamma_{max}}(\cdot)$ denotes the Cumulative Distribution Function (CDF) of $\gamma_{max}$, where $\gamma_{max}=\bar{\gamma}\left({\textstyle\sum_{n=1}^{N_{eff}}\eta_{n}}\right)^{2}$, where we denote $\eta_{n}=\alpha_{n}\beta_{n}$. For simplicity, we adopt the description same as [3], which allows us to obtain $F_{\gamma_{max}}=\frac{\Gamma\left(\Lambda N_{eff},\sqrt{\frac{\gamma_{th}}{\bigtriangleup^{2}\bar{\gamma}}}\right)}{\Gamma\left(\Lambda N_{eff}\right)}$, where $\Gamma\left(\cdot,\cdot\right)$ is the incomplete Gamma function, while $\Gamma\left(\cdot\right)$ is the Gamma function, parameters are set as $\Lambda=\frac{\pi^{2}}{16-\pi^{2}}$, and $\bigtriangleup=\frac{\left(16-\pi^{2}\right)}{2\pi}$ following [4].
Figure 2: Outage probability $P_{out}$ performance comparison where HPBW $\phi_{0}=13.8^{\circ}$, $y_{s}=15$m, and $\gamma_{th}=20$ dB: (a) $P_{out}$ versus transmit power $P_{t}$, (b) $P_{out}$ versus elements number NN, where transmit power $P_{t}=-20$ dBm.
@@ III Performance Analysis
We analyze the performance in the near field and assume that the elements within the MA-RIS configuration shift horizontally, transitioning from the panel’s left edge to its right edge. The positions of Tx and Rx are set as $(x_{b},y_{b},z_{b})=(0,0,3)$ m and $(x_{u},y_{u},z_{u})=(10,0,0)$ m, The RIS panel is on the xy-plane, so $z_{c}=0$ m, x-axis’s center position is $x_{c}=5$ m, $y_{c}$ changes along with the RIS panel’s position during each iteration in y-axis, the length of RIS panel is set as l=1l=1m, and the number of elements is $N=21$. The antenna movement velocity is $v=0.1$ m/s, with the traversal time thus amounting to $T=10$s. Following [11], we set the carrier frequency as $f=4.25$ GHz, and the noise variance as $\sigma^{2}=-45$ dBm.
Figure 3: Outage probability $P_{out}$ and Instantaneous SNR $\gamma$ versus y-axis position $y_{s}$ for HPBW $\phi_{0}=13.8^{\circ}$, and $\gamma_{th}=20$ dB.
We assume that all the $N$ MA elements move together with speed $v$ along the RIS panel of length $l$ m within a duration time of $T$ s. The RIS panel is symmetrically set at $\textbf{r}_{c}$ on the xy-plane while it also moves vertically along y-axis. So we can write the offset from the RIS panel’s center $x_{c}$ at each time instant as $d_{t,\mathrm{off}}=(t-1)v-l/2$, where $t\in\left\{1,2,\cdots,T\right\}$. So we could get the center element’s position at each time instant $x_{t,c}=x_{c}+d_{i,\mathrm{off}}$, and in turn each MA element’s x-axis position $x^{r}_{n}=x_{t,c}+(n-1)*d_{sm},n\in\left\{1,2,\cdots,N\right\}$, where $d_{sm}$ is the element spacing between MAs.
@@ A Transmit Power and Number of Elements : Effect on Outage Probability
As shown in Fig. 2LABEL:sub@OutvsPt, the outage probability decreases with an increase in transmit power for all configurations. In 2D scenario with 21 elements and -20 dBm transmit power, the MA-RIS achieves around 24% better performance than FPA-RIS in terms of outage probability. In a 2D scenario,
the outage probability sharply decreases with increasing transmit power due to better control over signal propagation and interference mitigation. This is because the additional spatial dimension enables more precise optimization of signal paths, resulting in improved channel conditions. The impact of number of elements on outage probability $P_{out}$ is shown in Fig. 2LABEL:sub@OutvsN, with the transmit power $P_{t}$ fixed at -20 dBm. Due to the single-path scenario we consider, there is no significant difference between the 1D and 2D configurations for both the MA-RIS and FPA-RIS settings, due to the fact the outage probability is jointly determined by $\bar{\gamma}$ and $N_{eff}$ as $P_{out}\to\left(\frac{\bar{\gamma}\triangle}{\gamma_{th}}\Gamma\left(\Lambda N_{eff}\right)^{-\frac{2}{\Lambda N_{eff}}}\right)^{-\frac{\Lambda N_{eff}}{2}}$ [4]. As we can see from Fig. 3, when the MA-RIS panel reaches the highest heights, $P_{out}$ and $\gamma$ gradually approach a common value for the 1D and 2D cases, matching what is shown in Fig. 2LABEL:sub@OutvsN. In particular, the results show that MA-RIS uses fewer elements than FPA-RIS to achieve the same level of performance, for example for an outage probability of $10^{-70}$, MA-RIS uses 25% fewer elements than FPA-RIS.
The maximum instantaneous SNR $\gamma$, was determined at positions $y_{s}=4.5,4,6,4$ m, yielding instantaneous SNR values of 16.8, 17.2, 14.8, and 15.3 dB for MA-1D, MA-2D, FPA-1D, and FPA-2D, respectively.
These variations in the optimal positions at $y_{s}$ are attributed to differences in geometric compactness. Given that channel state information is assumed to be known, the average SNR, $\bar{\gamma}$, follows a similar trend to the instantaneous SNR, $\gamma$. Consequently, the $P_{out}$ is inversely proportional to the instantaneous SNR and is influenced by both $\gamma$ and the effective number of elements $N_{eff}$ based on (12). As $N_{eff}$ increases within a relatively lower height range, $P_{out}$ initially decreases; however, once $N_{eff}$ reaches its maximum, $P_{out}$ begins to increase as $\gamma$ decreases.
These promising results are the consequence of MA-enabled dynamic reconfiguration of antenna elements, allowing for adaptive adjustment of the antenna pattern and phase distribution. This flexibility can be leveraged to optimize the signal path and mitigate multipath fading, leading to enhanced communication reliability and lower outage probabilities compared to FPA-RIS, which instead have a limited capability to adapt to changing channel conditions.
@@ B Effective Number of MA-RIS
We assume that both the MA-RIS and FPA-RIS panels move vertically along the y-axis, denoted as $y_{s}=y_{r}$ in Fig. 4. They move from an initial position of $y_{s}=$1 m to the maximum height of $y_{s}=15$ m in increments of 2 m in xy-plane. In order to compare the performance of FPA-RIS and MA-RIS over the same overall panel length, the antenna spacing for FPA-RIS is set as $d_{sf}=\lambda/2$, while the antenna spacing for MA-RIS is $(vt-(N-1)d_{sf})/(N-1)$.
We can observe that for both 1D and 2D configurations for the MA-RIS, the effective antenna number $N_{eff}$ reaches its maximum value for lower values of $y_{s}$ as compared to FPA-RIS. This indicates that the ability to adjust the position of the RIS panel allows one to optimize the reflective path and thus the effective utilization of more antenna elements. We also notice that in 2D configurations, the RIS can quickly reach its maximum $\textit{N}_{eff}$ as it begins to move from the initial position along the y-axis. This implies that, the 2D RIS has a higher initial capacity for effective signal reflection due to its expanded interaction field, reaching peak performance with minimal movement. Instead, for 1D configurations, the effectiveness of the RIS in terms of $\textit{N}_{eff}$ increases more significantly as one moves to higher positions along the y-axis. In fact, in 1D setups the initial positions may not be optimal for capturing and redirecting signals effectively, while as the RIS moves to higher y-axis positions, it can encounter angles and positions where the linear array of antennas becomes more aligned with the direction of the incident signals, thereby increasing the effective number of antennas that can contribute to signal reflection.
Figure 4: Comparisons of effective number of elements in MA-RIS and FPA-RIS versus y-axis position $y_{s}$ where HPBW $\phi_{0}=13.8^{\circ}$, the center position of RIS panel is $x_{c}=5$ m along x-axis.
Figure 5: Instantaneous SNR $\gamma$ versus transmit power $P_{t}$, with noise variance $\sigma^{2}=-45$ dBm, and y-axis position $y_{s}=15$ m.
@@ C Instantaneous SNR
In Fig. 5, we compare instantaneous SNR versus the transmit power $P_{t}$. We can see for example that when the transmit power $P_{t}$ is 0 dBm, then the MA-RIS’s instantaneous SNR is about 3 dB and 2.5 dB higher than FPA-RIS in 1D and 2D setups, respectively. This is due to the fact that MA-RIS systems have the inherent advantage of being able to adapt their position and orientation, in response to the environment and the direction of the incoming signal. This adaptability allows them to optimize the reflection path, enhancing the signal’s strength at the receiver and thereby improving the SNR. The superior performance of 2D is leaded by the geometry compactness. 2D configurations can enable more sophisticated beamforming strategies, leveraging the additional spatial dimension to focus the energy more precisely towards the receiver. Moreover, 2D RIS configurations can accommodate a wider range of AoA and AoD values for the signal. This capability enables them to maintain high SNR levels over a variety of transmission and reception scenarios, unlike 1D RIS, which is more limited in this respect.
@@ IV Conclusion
Our work highlights the importance of considering the following factors with regard to antenna elements: the element geometry distribution, i.e., 1D or 2D, and each element’s position in RIS panel, as well as the position of the RIS panel. Those factors are deeply connected with the angles of incidence and interaction with reflection signals, which contribute the signal strength and coverage. This could prove particularly beneficial in dynamic or large-scale environments, e.g., V2X and Massive MIMO communication systems, since the flexible adjustment of the antenna positions could provide an important additional degree of freedom to capture the change in environment characteristics, resulting in enhanced coverage, and improved signal quality. Despite our work highlight the advantages of MA-RIS, the limitations inherent in our current model set the stage for our future efforts to refine and expand its capabilities. The model’s idealized assumptions, such as a fixed single-user position and a single-path scenario, led to a notably low outage probability. Our future work aims to address these limitations by incorporating more realistic multi-user mobility scenarios and multipath propagation conditions, enhancing both the model’s applicability and accuracy.